in a scene in an action movie, a stuntman jumps from the top of one building to the top of another building 4.0 m away. after a running start, he leaps at a velocity of 5.0 m/s at an angle of 15* with respect to the flat roof. will he make it to the other roof, which is 2.5 m shorter than the building he jumps from?


1
Expert's answer
2021-01-07T10:55:45-0500
"t=\frac{x}{v\cos{15}}=\frac{4}{5\cos{15}}=0.833\ s"

"y=vt\sin{15}-0.5gt^2\\y=(5)(0.833)\sin{15}-0.5(9.8)(0.833)^2=-2.35\ m"

So, he will make it to the other roof as 2.35 m is less than 2.50 m. He is able to cross the cliff without any injury.